C=260+0.75y I=320 G=300


1) Determine the equilibrium level of income using expenditure and injection leakage.


2) calculate the value of multiplier of economy


3) Calculate the equilibrium level of income when there is a decrease in investment from 320 to 300 using expenditure and injection leakage approach


1
Expert's answer
2021-11-14T17:41:02-0500

Solution:

1.). At equilibrium level of income: AS = AD

Y = C + I + G

Y = 260 + 0.75Y + 320 + 300

Y – 0.75Y = 260 + 320 + 300

0.25Y = 880

Y = 3,520

The equilibrium level of income = 3,520

 

2.). Multiplier = "\frac{1}{1 - MPC}"

MPC = 0.75

Multiplier = "\frac{1}{1 - 0.75} = \frac{1}{0.25} = 4"


The value of the multiplier = 4

 

3.). New equilibrium level of income:

Y = C + I + G

I = 300

Y = 260 + 0.75Y + 300 + 300

Y – 0.75Y = 260 + 300 + 300

0.25Y = 860

Y = 3,440

New equilibrium level of income = 3,440